OpenStudy (anonymous):

Solve log12 (x + 1) - log12 4 = log12 16 Question 4 options: 1) x = 3 2) x = 12 3) x = 63 4) x = 64

OpenStudy (anonymous):

same way dude...as previous

OpenStudy (anonymous):

log12((x+1)/4))=log12(16)

OpenStudy (anonymous):

got it??

OpenStudy (anonymous):

idk how you got that but yea sure

OpenStudy (anonymous):

Ok to start with logarithms you must apply all the logarithm properties. The one you are going to use for this exercise is as follows: \[\log_{A}B-\log_AC=\log_A[B/C]\]the second property that you must use is the following:\[A ^{\log_AB}=B\]. Now that you these properties, lets use them:\[\log_{12}(x+1)-\log_{12}4=\log_{12}[(x+1)/4]\] and with the second property we have that \[\log_{12}[(x+1)/4]=\log_{12}16\] which implies that\[12^{\log_{12}[(x+1)/4]}=12^{\log_{12}16}\] which implies that\[[(x+1)/4]=16\] which clearly implies that\[x+1=64\rightarrow x=63\]. Peace, Love and Happiness from Puerto Rico